ROTATIONAL TEMPERATURE MEASUREMENTS OF AN OPTICALLY PUMPED, VIBRATIONALLY EXCITED CARBON MONOXIDE-ARGON PLASMA USING SINGLE PHOTON LASER INDUCED FLUO-RESCENCE OF THE (v′′ = 20)X^{1}Σ^{+} ←(v′ = 2)D^{1}Σ^{+} BAND

\underline{R. J. LEIWEKE} and W. R. LEMPERT, Nonequilibrium Thermodynamics Laboratory, Department of Mechanical Engineering and Department of Chemistry, The Ohio State University, Columbus, OH 43210.